# InfinitePOD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

### v2

@@ Abstract.
We propose InfinitePOD, a novel transceiver-centric HBD architecture that unifies connectivity and dynamic switching at the transceiver level using Optical Circuit Switching (OCS). By embedding OCS within each transceiver, InfinitePOD achieves reconfigurable point-to-multipoint connectivity, allowing the topology to adapt into variable-size rings. This design provides: i) datacenter-wide scalability without cost explosion; ii) fault resilience by isolating failures to a single node, and iii) full bandwidth utilization for fault-free GPUs. Key innovations include a Silicon Photonic (SiPh) based low-cost OCS transceiver (OCSTrx), a reconfigurable k-hop ring topology co-designed with intra-/inter-node communication, and an HBD-DCN orchestration algorithm maximizing GPU utilization while minimizing cross-ToR datacenter network traffic. The evaluation demonstrates that InfinitePOD achieves 31% of the cost of NVL-72, near-zero GPU waste ratio (over one order of magnitude lower than NVL-72 and TPUv4), near-zero cross-ToR traffic when node fault ratios under 7%, and improves Model FLOPs Utilization by 3.37x compared to NVIDIA DGX (8 GPUs per Node).
@@ 1. Introduction
Based on these principles, we propose InfinitePOD, a scalable and fault-resilient HBD architecture designed for optimizing TP communication. Our key insight is unifying connectivity and dynamic switching at the transceiver level using OCS. By embedding OCS in each transceiver, we achieve reconfigurable point-to-multipoint connectivity. This marks a departure from traditional designs, where transceivers support only point-to-point connections and rely on high-radix switches for routing. We call this new design transceiver-centric HBD architecture. This transceiver-centric architecture offers two key benefits: i) It enables the flexible construction of arbitrarily large ring topologies by intra-node loopback mechanism. This can support optimal TP group sizes for different models, while effectively minimizing resource fragmentation; ii) When one node fails, its neighboring transceivers dynamically reconfigure connections to reroute traffic, significantly reducing the fault explosion radius and improving system resilience.
@@
To the best of our knowledge, InfinitePOD is the first HBD design capable of scaling to datacenter scale while avoiding cost explosion and increased failure-induced waste. We evaluated InfinitePOD with the real 160-day fault trace from our 3K GPU cluster [^2] Details in Appendix §A. We will open-source the trace later.. When executing TP32 jobs with the trace, it demonstrates 0.44% GPU waste ratio - 22x and 23x lower than NVL-72 (9.67%) and TPUv4 (10.30%). It achieves 3.24x and 1.59x cost reductions compared to NVIDIA NVL-72 and Google TPUv4 respectively. Through the orchestration algorithm, it maintains near-zero cross-ToR traffic under 7% node failure rates. Its dynamic ring formation capability enables 3.37x higher MFU than NVIDIA DGX systems (Platform, 2024) (8 GPUs/node).
@@ 2.1. LLM Training in AI DC
InfinitePOD Transceiver-centric High Ring-Allreduce Node-level Low Few
@@ 2.2. Limitations of Existing HBDs
Figure 2. InfinitePOD overview.
@@ 3. Design Overview
In this section, we first present our new HBD architecture InfinitePOD guided by the design principles outlined above. We then provide an overview of its key components.
@@
We realize this design with InfinitePOD, which has three key components as shown in Figure 2.
@@
4. InfinitePOD Design
This section first introduces the innovative design of OCS transceivers (OCSTrx) based on Silicon Photonics (SiPh) chips (§4.1), a key enabler for InfinitePOD, providing both cost efficiency and reconfigurability. Next, we present the DC-scale InfinitePOD topology design (§4.2) based on OCSTrx. Finally, we outline the HBD-DCN orchestration algorithm (§4.3), designed to optimize communication efficiency for training jobs.
@@ 4.1. SiPh-based OCS transceiver (OCSTrx)
4.2. InfinitePOD Topology
In this section, we present the InfinitePOD topology design (Figure 2) integrating OCSTrx that allows all GPUs within datacenter to be connected in a reconfigurable K-Hop Ring topology, while supporting dynamic ring construction and high fault tolerance.
@@ 4.3. HBD-DCN Orchestration Algorithm
InfinitePOD is designed to work with arbitrary DCN, including Rail-Optimized (NVIDIA, 2021; Qian et al., 2024) and Fat-Tree (Al-Fares et al., 2008). This section co-optimizes communication performance for both HBD and DCN in InfinitePOD.
Problem Statement. In InfinitePOD, GPUs communicate without routing traffic, preventing congestion at any scale. In contrast, DCNs experience inevitable congestion, leading to performance degradation. To mitigate this, we leverage traffic locality to orchestrate nodes, minimizing cross-ToR traffic. Given a job $J$ requiring $N$ nodes from an available pool of $M$ ($M\geq N$), we must select and order $N$ nodes to satisfy two requirements: (1) nodes in the same TP group should communicate via InfinitePOD, and (2) other parallel traffic should minimize congestion. Ideally, communication remains within the same ToR, confining congestion to switch-to-node links.
@@
A naive approach is sorting nodes based on deployment order in InfinitePOD, fulfilling the first requirement but not the second. As shown in Figure 6(a), this method places ($N_{1}$, $N_{2}$) in the same TP group and ($N_{1}$, $N_{3}$) in the same DP group, forcing DP traffic across ToRs. A better scheme (Figure 6(b)) eliminates cross-ToR traffic and congestion. However, considering failures and multiple parallel dimensions complicates orchestration, necessitating an efficient method.
Our key insight is to arrange nodes in InfinitePOD based on DCN traffic locality, prioritizing appropriate network distances over minimal ones. For example, in Figure 6(b), $N_{1}$’s InfinitePOD neighbor is $N_{3}$, despite a 3-hop network distance in DCN. We propose a two-phase solution: (1) a deployment phase defining physical connections in DCN and InfinitePOD, and (2) a runtime phase using an algorithm to orchestrate nodes for arbitrary-scale jobs.
@@
Deployment Phase. Figure 7 shows node deployment in HBD and DCN. InfinitePOD connects nodes at a network distance of 3 (i.e., cross-ToR). In a DCN with $r$ nodes per ToR, node $N_{n}$ connects to $N_{n\pm r}$ as main links, while backup links connect to $N_{n\pm 2r}$. For $1<n\leq r$, $N_{n}$ connects to $N_{D+n-r-1}$, where $D$ is the total node count (e.g., $N_{3}$ connects to $N_{14}$). Additionally, $N_{1}$ may link to the last node, forming a ring.
Runtime Phase. Without considering DCN topology, InfinitePOD orchestrates nodes in three steps: (1) identifying cluster faults and modeling healthy nodes as a graph, (2) using Depth-First Search to find connected components, and (3) sequentially placing TP groups within these components. Due to InfinitePOD’s topology, each TP group forms a ring.
@@ 6. Large-Scale Simulation
We begin by outlining the experimental methodology and setup (§6.1). Next, we assess fault tolerance across different HBD architectures (§6.2), followed by end-to-end simulations to evaluate training performance under varying parallelism and GPU resource allocations (§6.3). We then examine the improvements in communication efficiency achieved by our orchestration algorithm (§6.4). Finally, we present a comparative cost and power analysis of different HBD architectures (§6.5). The simulations demonstrate that InfinitePOD outperforms other architectures across all metrics.
@@ 6.1. Methodology and Setup
An in-house simulator dedicated for LLM training is used to evaluate InfinitePOD comprehensively. The simulator supports end-to-end simulations of both model training performance and hardware faults, with the HBD-DCN orchestration algorithm seamlessly integrated into the system.
@@
Parallelism strategy. Since InfinitePOD is primarily designed for TP, the key variable is the TP size. TP-8, TP-16, TP-32, and TP-64 are tested to evaluate the fault resilience of various HBD architectures (§6.2). Additionally, other parallelism strategies, such as PP and DP, are used to simulate cross-ToR traffic and evaluate the orchestration algorithm (§6.4).
Fault patterns. The fault trace used in the simulation was collected from an 8-GPU node cluster with approximately 3,000 GPUs over a span of 160 days. On average, the ratio of faulty 8-GPU nodes is $3.83\%$, with the P99 value as $7.22\%$, more details in Appendix §A. In some simulations, fault traces generated based on this trace statistics are also derived.
@@ 6.2. HBD Fault Resilience
In these scenarios, InfinitePOD ($K=3$) achieves near-zero GPU waste ratio, and outperforming all other architectures. Especially, the waste ratio for InfinitePOD ($K=2$) remains almost identical to InfinitePOD ($K=3$), allowing one bundle of OCSTrx to be saved for clusters with low fault rates. NVL-36 and NVL-72 typically experience an 11% waste ratio for TP sizes of 16 or larger, as $1/9$ of GPUs are reserved for redundant backups. NVL-576 has less fragmentation, benefiting from its larger size. TPUv4 performs well at low fault ratios and small TP sizes, but significantly degrades with larger TP sizes due to its coarse $4^{3}$ cube-based resource management, which amplifies the fault explosion radius. To sum up, InfinitePOD demonstrates the strongest fault resilience among all architectures.
Maximum job supported. In fixed-size clusters, large job must pause when the available GPUs drop below the required count. Faced with same fault rate, cluster with lower GPU waste ratio can support larger job scales. Figure 11 shows the maximum job scale supported for various HBD architectures cluster with 2880-GPU, simulated with the fault traces normalized for 4-GPU nodes. InfinitePOD ($K=2$ or $K=3$) and NVL-576 lead in performance, and SiP-Ring exhibits declining efficiency as TP size increases.
Figure 11. Maximal job scale supported by 2880 GPUs.
Job fault-waiting time. Large job must wait for the repairing when GPU availability falls below the required threshold. This simulations assume the average recovery time in the fault trace as a fixed repair duration. The total wasted time during 160 days is evaluated (Figure 12). For smaller TP sizes (TP-8/TP-16), NVL-36/NVL-72 exhibit the weakest resilience due to their 11% backup overhead. For larger TP sizes (TP-32/TP-64), SiP-Ring and TPUv4 perform worst.
@@ 6.3. Training Performance
This section analyzes the training performance of two representative large models, LLama 3.1-405B (Llama Team, AI @ Meta, 2024) and GPT-MoE (configuration detailed in Appendix §B), under various GPU resource configurations and parallelism strategies. The simulation results validate the practical applicability of the InfinitePOD architecture. In simulations, we model practical TP and EP behaviors: For TP, increasing parallelism splits GEMMs into smaller, less efficient tasks, reducing hardware efficiency (NVIDIA, 2023); for EP, we practically set expert imbalance coefficient at 20%.
@@ 6.4. Communication Efficiency
This section examines the impact of orchestration algorithms on DCN communication efficiency. Experiments were performed on a Fat-Tree architecture, like the setup in  (Gangidi et al., 2024b). As shown in Figure 13(a), the algorithm is not sensitive to cluster size. Therefore, the evaluation is based on TP-32 operations on InfinitePOD with 8192 GPUs.
@@ 6.5. Cost and Power Analysis
The breakdown analysis of each architecture is provided in the Appendix §F. Based on this, the cost and power consumption are normalized according to GPU count and per-GPU bandwidth. As depicted in Table 6, InfinitePOD exhibits the lowest interconnect cost per GPU per GBps. Under the $K=2$ configuration, its cost is only 62.84% of Google TPUv4 and 30.86% of the NVIDIA GB200 NVL-36/72, with minimal power consumption. This efficiency is primarily attributed to the avoidance of centralized switches. TPUv4 ranks second in interconnect cost and lowest in power consumption, achieved by reducing optical module use and per-port OCS costs. The NVL series has higher interconnect costs and power consumption due to its fully-connected topology and high-cost NVLink Switches. Notably, NVL-576 incurs the highest cost and power consumption due to its multilayer nonconvergent topology, which increases optical module expenses and requires more NVLink Switches.
@@
Simulations on a 11,520-GPU cluster using the TP-32 configuration evaluate GPU availability under varying fault ratios across different architectures. The variation in aggregate cost for different HBD architectures under varying node fault ratios is illustrated in Figure 13(d). InfinitePOD consistently exhibits the lowest aggregate cost. Furthermore, when the fault ratio is below 12.1%, the aggregate cost of InfinitePOD ($K=2$) is less than that of InfinitePOD ($K=3$), suggesting that ($K=2$) is the optimal design for most scenarios.
@@ 7. Discussion
AllToAll communication. Ring topology in InfinitePOD struggles with AllToAll communication (e.g., EP), exhibiting poor performance at $O(p^{2})$, where $p$ is the group size. This can be improved by linking backup lines to nodes indexed at $n\pm 2^{i}$ instead of $n\pm i$ and Applying the Binary Exchange algorithm, and reduce time complexity to $O(p\log_{2}p)$. During the algorithm, OCSTrx need to connect to different GPUs with runtime switching, since OCSTrx switches in under $1ms$, reconfiguration can be overlapped with computation. For $K=2$ InfinitePOD designs, performance matches the ideal Bruck algorithm (Bruck et al., 1997) when $p<8$. However, this design introduces complexities in construction, failover, and orchestration, and necessitates GPU routing capabilities. So it is not applied.
Simulation Scale. Simulations using real 3,000 GPU fault traces, as detailed in §6.2, were conducted on a cluster comprising 2880 GPUs. The reason is that the simulation’s GPU count must be less than the total GPUs in the fault trace, and 2880 is the largest number divisible by 576 and less than 3,000. This configuration allows the entire cluster to be divided into five NVL-576 units for the simulation. Larger scales are simulated in other scenarios.
Multi-dimension parallelism. InfinitePOD is optimized for single-dimension parallelism. To support multi dimensional communication, two approaches are viable. 1) Independent Interconnects: Each OCSTrx bundle includes multiple OCSTrx units (e.g., 4 or 8), then link each of the unit to a separate inter-host topology. This isolates parallel dimensions but results in fixed bandwidth per dimension, leading to inefficiencies. 2) Time-Division multi-dimension: Main and backup lines of OCSTrx can be used to form separated inter-host topology, rapidly switching between them can support multi-dimension parallelism. However, this introduces complexity in managing multi-dimensional overlap and reduces the fault tolerance of InfinitePOD.
Single-Job vs. Multi-Job. Existing studies explore multi-job scheduling in GPU clusters (Weng et al., 2022; Li et al., 2023b). Deploying certain small jobs, such as inference tasks, can mitigate GPU fragmentation. However, given the shortage of GPU in LLM training, any idle GPU—whether repurposed for small job or not—is undesirable. Thus, InfinitePOD prioritizes single-job execution for simplicity.
OCS vs. EPS. OCSTrx enables multi-path selection, a feature also achievable with Electronic Packet Switching (EPS). For example, inter-host topology of InfinitePOD can be implemented using UBB 2.0-based servers by adding external optical interfaces to switches in server. However, this would require twice optical modules and numerous high-throughput switching chips for the entire system, significantly increasing cost and power consumption compared to OCSTrx.

### v3

# InfiniteHBD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

@@ Abstract.
We propose InfiniteHBD, a novel transceiver-centric HBD architecture that unifies connectivity and dynamic switching at the transceiver level using Optical Circuit Switching (OCS). By embedding OCS within each transceiver, InfiniteHBD achieves reconfigurable point-to-multipoint connectivity, allowing the topology to adapt into variable-size rings. This design provides: i) datacenter-wide scalability without cost explosion; ii) fault resilience by isolating failures to a single node, and iii) full bandwidth utilization for fault-free GPUs. Key innovations include a Silicon Photonic (SiPh) based low-cost OCS transceiver (OCSTrx), a reconfigurable k-hop ring topology co-designed with intra-/inter-node communication, and an HBD-DCN orchestration algorithm maximizing GPU utilization while minimizing cross-ToR datacenter network traffic. The evaluation demonstrates that InfiniteHBD achieves 31% of the cost of NVL-72, near-zero GPU waste ratio (over one order of magnitude lower than NVL-72 and TPUv4), near-zero cross-ToR traffic when node fault ratios under 7%, and improves Model FLOPs Utilization by 3.37x compared to NVIDIA DGX (8 GPUs per Node).
@@ 1. Introduction
Based on these principles, we propose InfiniteHBD, a scalable and fault-resilient HBD architecture designed for optimizing TP communication. Our key insight is unifying connectivity and dynamic switching at the transceiver level using OCS. By embedding OCS in each transceiver, we achieve reconfigurable point-to-multipoint connectivity. This marks a departure from traditional designs, where transceivers support only point-to-point connections and rely on high-radix switches for routing. We call this new design transceiver-centric HBD architecture. This transceiver-centric architecture offers two key benefits: i) It enables the flexible construction of arbitrarily large ring topologies by intra-node loopback mechanism. This can support optimal TP group sizes for different models, while effectively minimizing resource fragmentation; ii) When one node fails, its neighboring transceivers dynamically reconfigure connections to reroute traffic, significantly reducing the fault explosion radius and improving system resilience.
@@
To the best of our knowledge, InfiniteHBD is the first HBD design capable of scaling to datacenter scale while avoiding cost explosion and increased failure-induced waste. We evaluated InfiniteHBD with the real 348-day fault trace from our 10K GPU cluster [^2] Details in Appendix §A. We have open-sourced this trace at https://gitlab.basemind.com/sys/infra/superpod/infinitehbd-trace. When executing TP32 jobs with the trace, it demonstrates 0.15% GPU waste ratio - 69x and 45x lower than NVL-72 (10.31%) and TPUv4 (6.76%). It achieves 3.24x and 1.59x cost reductions compared to NVIDIA NVL-72 and Google TPUv4 respectively. Through the orchestration algorithm, it maintains near-zero cross-ToR traffic under 7% node failure rates. Its dynamic ring formation capability enables 3.37x higher MFU than NVIDIA DGX systems (Platform, 2024) (8 GPUs/node).
@@ 2.1. LLM Training in AI DC
InfiniteHBD Transceiver-centric High Ring-Allreduce Node-level Low Few
@@ 2.2. Limitations of Existing HBDs
Figure 2. InfiniteHBD overview.
@@ 3. Design Overview
In this section, we first present our new HBD architecture InfiniteHBD guided by the design principles outlined above. We then provide an overview of its key components.
@@
We realize this design with InfiniteHBD, which has three key components as shown in Figure 2.
@@
4. InfiniteHBD Design
This section first introduces the innovative design of OCS transceivers (OCSTrx) based on Silicon Photonics (SiPh) chips (§4.1), a key enabler for InfiniteHBD, providing both cost efficiency and reconfigurability. Next, we present the DC-scale InfiniteHBD topology design (§4.2) based on OCSTrx. Finally, we outline the HBD-DCN orchestration algorithm (§4.3), designed to optimize communication efficiency for training jobs.
@@ 4.1. SiPh-based OCS transceiver (OCSTrx)
4.2. InfiniteHBD Topology
In this section, we present the InfiniteHBD topology design (Figure 2) integrating OCSTrx that allows all GPUs within datacenter to be connected in a reconfigurable K-Hop Ring topology, while supporting dynamic ring construction and high fault tolerance.
@@ 4.3. HBD-DCN Orchestration Algorithm
InfiniteHBD is designed to work with arbitrary DCN, including Rail-Optimized (NVIDIA, 2021; Qian et al., 2024) and Fat-Tree (Al-Fares et al., 2008). This section co-optimizes communication performance for both HBD and DCN in InfiniteHBD.
Problem Statement. In InfiniteHBD, GPUs communicate without routing traffic, preventing congestion at any scale. In contrast, DCNs experience inevitable congestion, leading to performance degradation. To mitigate this, we leverage traffic locality to orchestrate nodes, minimizing cross-ToR traffic. Given a job $J$ requiring $N$ nodes from an available pool of $M$ ($M\geq N$), we must select and order $N$ nodes to satisfy two requirements: (1) nodes in the same TP group should communicate via InfiniteHBD, and (2) other parallel traffic should minimize congestion. Ideally, communication remains within the same ToR, confining congestion to switch-to-node links.
@@
A naive approach is sorting nodes based on deployment order in InfiniteHBD, fulfilling the first requirement but not the second. As shown in Figure 6(a), this method places ($N_{1}$, $N_{2}$) in the same TP group and ($N_{1}$, $N_{3}$) in the same DP group, forcing DP traffic across ToRs. A better scheme (Figure 6(b)) eliminates cross-ToR traffic and congestion. However, considering failures and multiple parallel dimensions complicates orchestration, necessitating an efficient method.
Our key insight is to arrange nodes in InfiniteHBD based on DCN traffic locality, prioritizing appropriate network distances over minimal ones. For example, in Figure 6(b), $N_{1}$’s InfiniteHBD neighbor is $N_{3}$, despite a 3-hop network distance in DCN. We propose a two-phase solution: (1) a deployment phase defining physical connections in DCN and InfiniteHBD, and (2) a runtime phase using an algorithm to orchestrate nodes for arbitrary-scale jobs.
@@
Deployment Phase. Figure 7 shows node deployment in HBD and DCN. InfiniteHBD connects nodes at a network distance of 3 (i.e., cross-ToR). In a DCN with $r$ nodes per ToR, node $N_{n}$ connects to $N_{n\pm r}$ as main links, while backup links connect to $N_{n\pm 2r}$. For $1<n\leq r$, $N_{n}$ connects to $N_{D+n-r-1}$, where $D$ is the total node count (e.g., $N_{3}$ connects to $N_{14}$). Additionally, $N_{1}$ may link to the last node, forming a ring.
Runtime Phase. Without considering DCN topology, InfiniteHBD orchestrates nodes in three steps: (1) identifying cluster faults and modeling healthy nodes as a graph, (2) using Depth-First Search to find connected components, and (3) sequentially placing TP groups within these components. Due to InfiniteHBD’s topology, each TP group forms a ring.
@@ 6. Large-Scale Simulation
We begin by outlining the experimental methodology and setup (§6.1). Next, we assess fault tolerance across different HBD architectures (§6.2), followed by end-to-end simulations to evaluate training performance under varying parallelism and GPU resource allocations (§6.3). We then examine the improvements in communication efficiency achieved by our orchestration algorithm (§6.4). Finally, we present a comparative cost and power analysis of different HBD architectures (§6.5). The simulations demonstrate that InfiniteHBD outperforms other architectures across all metrics.
@@ 6.1. Methodology and Setup
An in-house simulator dedicated for LLM training is used to evaluate InfiniteHBD comprehensively. The simulator supports end-to-end simulations of both model training performance and hardware faults, with the HBD-DCN orchestration algorithm seamlessly integrated into the system.
@@
Parallelism strategy. Since InfiniteHBD is primarily designed for TP, the key variable is the TP size. TP-8, TP-16, TP-32, and TP-64 are tested to evaluate the fault resilience of various HBD architectures (§6.2). Additionally, other parallelism strategies, such as PP and DP, are used to simulate cross-ToR traffic and evaluate the orchestration algorithm (§6.4).
Fault patterns. The fault trace used in the simulation was collected from an 8-GPU node cluster with approximately 10K GPUs over a span of 348 days. On average, the ratio of faulty 8-GPU nodes is $2.66\%$, with the P99 value as $7.92\%$, more details in Appendix §A. In some simulations, fault traces generated based on this trace statistics are also derived.
@@ 6.2. HBD Fault Resilience
In these scenarios, InfiniteHBD ($K=3$) achieves near-zero GPU waste ratio, and outperforming all other architectures. Especially, the waste ratio for InfiniteHBD ($K=2$) remains almost identical to InfiniteHBD ($K=3$), allowing one bundle of OCSTrx to be saved for clusters with low fault rates. NVL-36 and NVL-72 typically experience an 11% waste ratio for TP sizes of 16 or larger, as $1/9$ of GPUs are reserved for redundant backups. NVL-576 has less fragmentation, benefiting from its larger size. TPUv4 performs well at low fault ratios and small TP sizes, but significantly degrades with larger TP sizes due to its coarse $4^{3}$ cube-based resource management, which amplifies the fault explosion radius. To sum up, InfiniteHBD demonstrates the strongest fault resilience among all architectures.
Maximum job supported. In fixed-size clusters, large job must pause when the available GPUs drop below the required count. Faced with same fault rate, cluster with lower GPU waste ratio can support larger job scales. Figure 11 shows the maximum job scale supported for various HBD architectures cluster with 9,792-GPU, simulated with the fault traces normalized for 4-GPU nodes. InfiniteHBD ($K=2$ or $K=3$) and NVL-576 lead in performance, and SiP-Ring exhibits declining efficiency as TP size increases.
Figure 11. Maximal job scale supported by 9,792 GPUs.
Job fault-waiting time. Large job must wait for the repairing when GPU availability falls below the required threshold. This simulations assume the average recovery time in the fault trace as a fixed repair duration. The total wasted time during 348 days is evaluated (Figure 12). For smaller TP sizes (TP-8/TP-16), NVL-36/NVL-72 exhibit the weakest resilience due to their 11% backup overhead. For larger TP sizes (TP-32/TP-64), SiP-Ring and TPUv4 perform worst.
@@ 6.3. Training Performance
This section analyzes the training performance of two representative large models, LLama 3.1-405B (Llama Team, AI @ Meta, 2024) and GPT-MoE (configuration detailed in Appendix §B), under various GPU resource configurations and parallelism strategies. The simulation results validate the practical applicability of the InfiniteHBD architecture. In simulations, we model practical TP and EP behaviors: For TP, increasing parallelism splits GEMMs into smaller, less efficient tasks, reducing hardware efficiency (NVIDIA, 2023); for EP, we practically set expert imbalance coefficient at 20%.
@@ 6.4. Communication Efficiency
This section examines the impact of orchestration algorithms on DCN communication efficiency. Experiments were performed on a Fat-Tree architecture, like the setup in  (Gangidi et al., 2024b). As shown in Figure 13(a), the algorithm is not sensitive to cluster size. Therefore, the evaluation is based on TP-32 operations on InfiniteHBD with 8192 GPUs.
@@ 6.5. Cost and Power Analysis
The breakdown analysis of each architecture is provided in the Appendix §F. Based on this, the cost and power consumption are normalized according to GPU count and per-GPU bandwidth. As depicted in Table 6, InfiniteHBD exhibits the lowest interconnect cost per GPU per GBps. Under the $K=2$ configuration, its cost is only 62.84% of Google TPUv4 and 30.86% of the NVIDIA GB200 NVL-36/72, with minimal power consumption. This efficiency is primarily attributed to the avoidance of centralized switches. TPUv4 ranks second in interconnect cost and lowest in power consumption, achieved by reducing optical module use and per-port OCS costs. The NVL series has higher interconnect costs and power consumption due to its fully-connected topology and high-cost NVLink Switches. Notably, NVL-576 incurs the highest cost and power consumption due to its multilayer nonconvergent topology, which increases optical module expenses and requires more NVLink Switches.
@@
Simulations on a 10K-GPU cluster using the TP-32 configuration evaluate GPU availability under varying fault ratios across different architectures. The variation in aggregate cost for different HBD architectures under varying node fault ratios is illustrated in Figure 13(d). InfiniteHBD consistently exhibits the lowest aggregate cost. Furthermore, when the fault ratio is below 12.1%, the aggregate cost of InfiniteHBD ($K=2$) is less than that of InfiniteHBD ($K=3$), suggesting that ($K=2$) is the optimal design for most scenarios.
@@ 7. Discussion
AllToAll communication. Ring topology in InfiniteHBD struggles with AllToAll communication (e.g., EP), exhibiting poor performance at $O(p^{2})$, where $p$ is the group size. This can be improved by linking backup lines to nodes indexed at $n\pm 2^{i}$ instead of $n\pm i$ and Applying the Binary Exchange algorithm, and reduce time complexity to $O(p\log_{2}p)$. During the algorithm, OCSTrx need to connect to different GPUs with runtime switching, since OCSTrx switches in under $1ms$, reconfiguration can be overlapped with computation. For $K=2$ InfiniteHBD designs, performance matches the ideal Bruck algorithm (Bruck et al., 1997) when $p<8$. However, this design introduces complexities in construction, failover, and orchestration, and necessitates GPU routing capabilities. So it is not applied.
Simulation Scale. Simulations using real 10,000 GPU fault traces, as detailed in §6.2, were conducted on a cluster comprising 9,792 GPUs. The reason is that the simulation’s GPU count must be less than the total GPUs in the fault trace, and 9,792 is the largest number divisible by 576 and less than 10,000. This configuration allows the entire cluster to be divided into 17 NVL-576 units for the simulation.
Multi-dimension parallelism. InfiniteHBD is optimized for single-dimension parallelism. To support multi dimensional communication, two approaches are viable. 1) Independent Interconnects: Each OCSTrx bundle includes multiple OCSTrx units (e.g., 4 or 8), then link each of the unit to a separate inter-host topology. This isolates parallel dimensions but results in fixed bandwidth per dimension, leading to inefficiencies. 2) Time-Division multi-dimension: Main and backup lines of OCSTrx can be used to form separated inter-host topology, rapidly switching between them can support multi-dimension parallelism. However, this introduces complexity in managing multi-dimensional overlap and reduces the fault tolerance of InfiniteHBD.
Single-Job vs. Multi-Job. Existing studies explore multi-job scheduling in GPU clusters (Weng et al., 2022; Li et al., 2023b). Deploying certain small jobs, such as inference tasks, can mitigate GPU fragmentation. However, given the shortage of GPU in LLM training, any idle GPU—whether repurposed for small job or not—is undesirable. Thus, InfiniteHBD prioritizes single-job execution for simplicity.
OCS vs. EPS. OCSTrx enables multi-path selection, a feature also achievable with Electronic Packet Switching (EPS). For example, inter-host topology of InfiniteHBD can be implemented using UBB 2.0-based servers by adding external optical interfaces to switches in server. However, this would require twice optical modules and numerous high-throughput switching chips for the entire system, significantly increasing cost and power consumption compared to OCSTrx.